# Holistic Tokenizer for Autoregressive Image Generation

### v1

@@ 5 Conclusion
We introduce Hita, a novel holistic-to-local tokenizer for AR image generation. Beyond image reconstruction, it emerges with new features such as zero-shot style transfer and zeros-shot image in-painting. Extensive experiments conducted on class-conditional image generation tasks illustrate it can accelerate converge speed during training and significantly improve the synthesis quality. Additionally, in-depth analysis also consolidates the effectiveness of the designed holistic tokenizer.
6 Acknowlegement
This work has been supported by the National Key R&D Program of China (Grant No. 2022YFB3608300), Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923, 17215025) Theme-based Research (Grant No. T45-701/22-R) and Shenzhen Science and Technology Innovation Commission (SGDX20220530111405040). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust. We are deeply grateful to Lufan Ma for the contribution in polishing up this paper.
@@ 6 Acknowlegement
In this subsection, we first compare Hita with the other vanilla tokenizers. Next, we depict the differences between VAR [47] in detail. Later, further discussion on token fusion modules are presented.
@@ A.1 Comparison with vanilla image tokenizers
Following VQGAN [12], we adopted ${\ell}_{2}$-normalization into codebook vectors, low codebook vector dimension, and a codebook size of 16,384. Compared with other counterparts like VQGAN [12] and ViT-VQGAN [54], the proposed tokenizer represents an image with fewer tokens (569 v.s. 1024), while achieving a better reconstruction quality with 100% utilization for both holistic and patch-level codebooks. Additionally, we observed that our approach achieves a better rIS 198.5 compared with the VQGAN proposed in LlamaGen [46]. rIS quantifies the KL-divergence between the original label distribution and the logit distribution of reconstructed images after softmax normalization [42]. In other words, rIS measures the semantic consistency between the reconstructed images and the original ones. A higher rIS confirms that our holistic tokenizer is more effective at preserving the semantic consistency of the reconstructed images.
@@
Table 6: Comparison with other image tokenizers. ${}^{\text{oim.}}$ indicates the tokenizer is trained on OpenImages [24]. $\mathcal{Q}_{H}$/$\mathcal{Q}_{P}$ denotes the codebook usage in holistic and patch-level quantizer, respectively.
@@ A.2 Comparison with VAR
As discussed in Sec.2, VAR’s multi-scale tokens can also be considered as a combination of semantic tokens and patch tokens. However, as depicted in Table. 7 we find removing the initial coarse-scale tokens seldom effect its reconstruction. Meanwhile, linear probing conducted on the cumulative coarse-scale tokens reveals poor semantic information.
@@
Table 7: Analysis on VAR’s multi-scale tokens. Acc indicates top-1 accuracy for linear probing estimation on the ImageNet [8].
Appendix B Further explanation on token fusion module
@@
As shown in Fig. 6, If a vanilla transformer is used to construct the token fusion module, due to the existence of the skip connection and the patch-level tokens contain enough information for image reconstruction, the patch-level tokens can directly flow through the skip connections into the decoder to reconstruct the image. Thus, the token fusion module can learn a trivial solution, which overlooks the holistic tokens and leads to holistic codebook collapse (Fig. 6a). Once the last $k$ holistic tokens participate in the image reconstruction, the information of the patch tokens flowing through the skip connection is incomplete. It needs to interact with the holistic tokens to obtain complete information for image reconstruction(Fig. 6b). Through this simple operation, we can not only emphasize the holistic tokens but also avoid its codebook collapse. To better align the token sequence with the nature of the AR generation model, here we adopt causal attention to build the token fusion module.
Appendix C More ablation study.
In this subsection, we elaborate more ablation studies on the design of our approach. Next, we conduct a quantitative analysis of the zero-shot in-painting performance of Hita.
@@ C.2 Other token fusion variants
As depicted in Sec 3.2.3, for simplicity, we choose the last $k$ holistic tokens combined with the patch-level tokens to reconstruct the image. Here, we refer to MAE [15] and Titok [59] and design 2 different variants to generate features for the first $k$ patch tokens: 1). Partial: A mask token combined with the positional embedding of the first $k$ patch tokens are used to generate their feature. In this scenario, the holistic tokens can be treated as condition; 2). Full: The patch tokens are completely removed, which is consistent with TiTok [59]. Then a mask token along with the positional embeddings is used to generate the features for all patch tokens. To achieve this, we initialize and train 2 different holistic tokenizers for 40 epochs, and then train 2 AR generation models – Hita-B and Hita-L based on those tokenizers with default training settings for 50 epochs. As listed in Table 9, the token fusion strategy proposed in Hita shows a better performance in both reconstruction and generation compared to the other variants. Thus, we choose to use the last ${k}$ holistic tokens combined with the patch-level tokens to reconstruct the image, by default.
@@
Table 9: Image reconstruction and AR generation. with other token fusion strategies.
@@
Table 10: Attention modules study in Hita. ‘-’ indicates the module is discarded from the tokenizer’s architecture.
C.4 Quantitative study on the in-painting quality.
@@
Evaluation metrics. To quantitatively estimate the semantic consistency of the generated images, we adopt top-1 accuracy and CLIP-score as our metrics, along with the generation metrics FID and IS. The top-1 accuracy is derived from image classification tasks, where we utilize a pre-trained ResNet-101 [14] on ImageNet [8] to classify the completed images. CLIP-score measures the similarity between the original image and the completed one. Specifically, we feed both original and in-painted images into the CLIP [36] model to extract image features and compute their cosine distance averaged across all samples. Larger top-1 accuracy and CLIP-score indicate a better semantic consistency is maintained in an AR generation model.
Observation and discussion. As shown in Fig. 7 and Fig. 9, the AR generation model trained with vanilla VQGAN [12] encounters difficulties in producing an image that maintains semantic coherence. In contrast, our approach can effectively produce visually consistent content to complete the given image part, maintaining a better overall coherence even with a significantly truncated prefix. Specifically, as observed in Fig. 7.a and Fig. 7.b, our method achieves higher classification accuracy and CLIP-score under various settings, with relatively stable fluctuations in all metrics across different configurations. In contrast, models trained with vanilla VQGAN exhibit more obvious performance variations, especially when only a small portion of the image is provided.
Beyond semantic coherency measurement, we also estimate the completed image quality with generation evaluation metrics, FID and IS. As shown in Fig. 7.c and Fig. 7.d, compared to the model trained with vanilla VQGAN[12], our approach achieves a better FID and IS, which also illustrates the holistic tokenizer can help the AR generation model generate the lower half of images robustly, maintaining strong semantic consistency.
Appendix D Limitation
Currently, Hita is trained on a limited dataset with vanilla optimization techniques, better performance could be achieved with more data and advanced learning objectives. Additionally, Hita can be seamlessly extended to perform text-conditional image generation, which is an ongoing direction of our research.
Appendix E More Visual Samples

### v4

@@ 5 Conclusion
Autoregressive (AR) models struggle with image generation because their unidirectional attention mechanisms restrict their ability to capture global context, resulting in lower generation quality. Hita addresses this by introducing a holistic-to-local tokenization scheme that uses learnable holistic tokens and local patch tokens to incoporate global information. Beyond image tokenization, Hita exhibits new features such as zero-shot style transfer and zeros-shot image in-painting. Experiments conducted on class-conditional image generation tasks illustrate that it accelerates convergence speed during training and significantly improves the synthesis quality. Besides, in-depth analysis also consolidates its effectiveness.
Acknowledgments
This work has been supported by the National Key R&D Program of China (Grant No. 2022YFB3608300), Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923, 17215025) Theme-based Research (Grant No. T45-701/22-R) and Shenzhen Science and Technology Innovation Commission (SGDX20220530111405040). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust. We are deeply grateful to Lufan Ma for the contribution in editing this paper.
@@ Acknowledgments
In this subsection, we first compare Hita with the other vanilla tokenizers. Next, we depict the differences between VAR [47] in detail. Next, we present further discussion on token fusion modules.
@@ A.1 Comparison with vanilla image tokenizers
Following VQGAN [12], we adopted ${\ell}_{2}$-normalization into codebook vectors, low codebook vector dimension, and a codebook size of 16,384. Compared with other counterparts like VQGAN [12] and ViT-VQGAN [54], the proposed tokenizer represents an image with fewer tokens (569 vs. 1024), while achieving a better reconstruction quality with 100% utilization for both holistic and patch-level codebooks. Additionally, we observed that our approach achieves a better rIS 198.5 compared with the VQGAN proposed in LlamaGen [46]. rIS quantifies the KL-divergence between the original label distribution and the logit distribution of reconstructed images after softmax normalization [42]. In other words, rIS measures the semantic consistency between the reconstructed images and the original ones. A higher rIS confirms that our holistic tokenizer is more effective at preserving the semantic consistency of the reconstructed images.
@@
Table 6: Comparison with other image tokenizers. ${}^{\text{oim.}}$ indicates training on OpenImages [24]. $\mathcal{Q}_{H}$/$\mathcal{Q}_{P}$ denote the codebook usage in holistic and patch-level quantizers, respectively.
@@ A.2 Comparison with VAR
As discussed in Sec.2, VAR’s multi-scale tokens can also be considered as a combination of semantic tokens and patch tokens. However, as depicted in Table. 7 we find removing the initial coarse-scale tokens seldom effects its reconstruction. Meanwhile, linear probing conducted on the cumulative coarse-scale tokens reveals poor semantic information.
@@
Table 7: Analysis of VAR’s multi-scale tokens. Acc indicates top-1 accuracy for linear probing estimation on the ImageNet [8].
Appendix B Explanation on Token Fusion Module
@@
As shown in Fig. 6, If a vanilla transformer is used to construct the token fusion module, due to the existence of the skip connection and the patch-level tokens contain enough information for image reconstruction, the patch-level tokens can directly flow through the skip connections into the decoder to reconstruct the image. Thus, the token fusion module can learn a trivial solution, which overlooks the holistic tokens and leads to holistic codebook collapse (Fig. 6a). Once the last $k$ holistic tokens participate in the image reconstruction, the information of the patch tokens flowing through the skip connection is incomplete. It needs to interact with the holistic tokens to obtain complete information for image reconstruction(Fig. 6b). This simple operation emphasizes the holistic tokens and avoids codebook collapse. To better align the token sequence with the nature of the AR generation model, here we adopt causal attention to build the token fusion module.
Appendix C More Ablation Studies.
We elaborate on further ablation studies on the design of our approach. Next, we quantitatively analyze Hita’s zero-shot inpainting performance.
@@ C.2 Other Token Fusion Variants
As depicted in Sec 3.2.3, for simplicity, we choose the last $k$ holistic tokens combined with the patch-level tokens to reconstruct the image. Here, we refer to MAE [15] and Titok [59] and design 2 different variants to generate features for the first $k$ patch tokens: 1). Partial: A mask token combined with the positional embedding of the first $k$ patch tokens are used to generate their feature. In this scenario, the holistic tokens can be treated as condition; 2). Full: The patch tokens are completely removed, which is consistent with TiTok [59]. Then, a mask token, along with the positional embeddings, generates the features for all patch tokens. To achieve this, we initialize and train 2 different holistic tokenizers for 40 epochs, and then train 2 AR generation models – Hita-B and Hita-L based on those tokenizers with default training settings for 50 epochs. As listed in Table 9, the token fusion strategy proposed in Hita shows a better performance in both reconstruction and generation compared to the other variants. Thus, we choose to use the last ${k}$ holistic tokens combined with the patch-level tokens to reconstruct the image, by default.
@@
Table 9: Image reconstruction and AR generation. with different token fusion strategies.
@@
Table 10: Attention modules analysis in Hita. ‘-’ indicates the module was removed from the tokenizer’s architecture.
C.4 Quantitative Study of In-painting Quality.
@@
Evaluation metrics. To quantitatively estimate the semantic consistency of the generated images, we adopt top-1 accuracy and CLIP-score as our metrics, along with the generation metrics FID and IS. The top-1 accuracy is derived from image classification tasks, where we utilize a pre-trained ResNet-101 [14] on ImageNet [8] to classify the completed images. CLIP-score measures the similarity between the original and the completed images. Specifically, we feed both original and in-painted images into the CLIP [36] model to extract image features and compute their cosine distance averaged across all samples. Higher top-1 accuracy and CLIP-scores indicate a better semantic consistency is maintained in an AR generation model.
Observation and discussion. As shown in Fig. 7 and Fig. 9, the AR generation model trained with vanilla VQGAN [12] encounters difficulties in producing an image that maintains semantic coherence. In contrast, our approach effectively produces visually consistent content to complete the given image part, maintaining a better overall coherence even with a significantly truncated prefix. Specifically, as observed in Fig. 7.a and Fig. 7.b, our method achieves higher classification accuracy and CLIP-score under various settings, with relatively stable fluctuations in all metrics across different configurations. In contrast, models trained with vanilla VQGAN exhibit more obvious performance variations, especially when only a small portion of the image is provided.
Beyond measuring semantic coherence, we also estimate the completed image quality using the generation evaluation metrics, FID and IS. As shown in Fig. 7.c and Fig. 7.d, compared to the model trained with vanilla VQGAN[12], our approach achieves a better FID and IS, which also illustrates the holistic tokenizer can help the AR generation model generate the lower half of images robustly, maintaining strong semantic consistency.
Appendix D Limitations
Currently, Hita is trained on a limited dataset using basic optimization techniques. Better performance could be achieved with more data and advanced learning objectives. Additionally, Hita can be seamlessly extended to perform text-conditional image generation, which is an ongoing direction of our research.
Appendix E More Visual Examples